# FP-Inconsistent

### v1

@@ 6 Inconsistency analysis
Based on the different takeaways in Section 5.2, we see that presenting certain values to different browser attributes helps evasive bots with evasion against anti-bot services. One way in which evasive bots could accomplish this would be to send requests from devices that would present the desired values for attributes. For example, evasive bots could send requests from a device containing 4 CPU cores to present a value of 4 for the hardwareConcurrency attribute. Alternatively, evasive bots could alter browser APIs and properties of their devices to present the desired values for different browser attributes. In this case, an evasive bot could alter the hardwareConcurrency attribute of the navigator browser object to present a desired value for the attribute, even though the device may have a different value for the number of CPU cores.
@@
In this section, we describe various inconsistencies observed in the values of different browser attributes among the requests received on our honey site. These inconsistencies provide evidence of bots manipulating browser APIs since such inconsistencies are extremely unlikely to occur when using real devices. We explore inconsistencies along a spatial and a temporal axis. Spatial inconsistencies constitute inconsistencies among browser attributes in a given request. Temporal inconsistencies constitute inconsistencies in browser attributes across requests from the same device. We use insights obtained from these inconsistencies to develop FP-Inconsistent, our semi-automated technique to generate inconsistency rules to detect evasive bots (described in Section 7).
@@
Figure 8: Bar plot showing the top 4 device types (inferred from the User-Agent) that have the highest probability of evasion against DataDome.
6.1 Spatial inconsistencies across browser attributes
Figure 8 shows the top 4 device types (inferred using the User-Agent property of the browser’s navigator object) that have the highest probability of evading DataDome based on the requests recorded on our honey site. From the figure, we see that iPhones have the highest probability of evasion (around 50%). We now look at the values of other browser attributes to answer if evasive bots sent requests from real iPhones or if they manipulated the navigator object on their browser to have their devices appear as iPhones. If bots are manipulating browser APIs and properties, we hypothesize that it would be difficult for them to ensure that all browser attributes convey consistent information. Thus, inconsistency across browser attributes indicates tampering of browser attributes which can be leveraged to improve bot detection since the browsers from real users are not likely to provide inconsistent values for attributes.
@@ 6.1 Spatial inconsistencies across browser attributes
To validate our hypothesis, we look into other attributes of requests have iPhones as their device type based on the User-Agent. Since we know that iPhones have a fixed set of screen resolutions (12 resolutions [41]), we inspect the spread of screen resolutions captured on requests from iPhones. Upon inspection, we find a total of 83 unique configurations from iPhones in the requests we received on our honey site, out of which 42 different resolutions were present among those requests from iPhones that evaded DataDome. This provides strong evidence to back our hypothesis that all bots that claim to use iPhones in their User-Agent do not originate from iPhones, since only 12 different resolutions are possible for iPhones in the real world.
Figure 9: Bar plot showing the top 10 screen resolutions among requests received from iPhones (when inferred using the User-Agent) that have the highest probability of evasion against DataDome. 9 out of these 10 resolutions do not exist in the real world indicating an inconsistency that can be leveraged to detect bots.
Figure 9 shows bar plots of the top 10 screen resolutions on iPhones that have the highest probability of evasion in the requests we received on our honey site. iPhones in the real world do not support 9 out of 10 of these screen resolutions, showing that the evasive bots that evaded DataDome had attributes in their browser configurations that were inconsistent with iPhones. In Section 7 we discuss our systematic, data-driven, semi-automatic approach to discover such inconsistencies to improve bot detection.
@@
Takeaway 6: While bots manipulate browser attributes to evade detection, it is difficult for them to ensure that all browser attributes are consistent with their manipulation. This provides opportunities to improve bot detection via inconsistencies. One particular value for a given attribute mapping to a large number of values for another attribute is a potential avenue to discover inconsistencies.
@@
Figure 10: Plots showing a heatmap of the geographical location of requests inferred using the timezone attribute of the navigator object and the IP address. Different regions lighting up in the two heatmaps indicate that while bots alter the navigator object or IP address or both to change their geographical location, they do not ensure that the location inferred using both is consistent.
6.2 Spatial inconsistencies across browser attributes and IP addresses
@@
Some bot services advertised sending traffic from specific geographic regions (USA, Mexico, France etc). Having this ability to send requests from specific regions suggests that the bot services are likely manipulating attributes that capture the geographical location of their devices. This manipulation introduces potential inconsistencies if the bot services did not ensure that all attributes point to the same region.
@@
Concretely, we analyzed requests sent from 4 different bot services who claimed to send requests from the United States, Canada, Europe and France respectively. We first used MaxMind’s GeoLite 2 database [42] to extract the geolocation from the IP address of the requests obtained from these services. We took a conservative approach when matching locations where we considered locations at the same UTC offset to be a match. For example, when analyzing requests from the vendor who advertised sending requests from France, we considered all requests whose geolocations mapped to any valid UTC offset that could overlap with France (such as Europe/Berlin) to also originate from France. With this approach, we observe that over 90% of requests from each of the 4 bot services originated from the advertised geographical location.
@@
However, we observed significant differences when repeating the same analysis using the browser’s timezone API [43] to infer location. We still used the same conservative approach of matching UTC offsets and merely replaced the geolocation inferred from the IP address with the timezone. We observed that only 76.52% of requests mapped to UTC offsets in Canada among the requests from the bot service that advertised traffic from Canada. More alarmingly, we observed that only 56% of requests mapped to UTC offsets in Europe among the requests from the bot service that advertised traffic from Europe. In contrast, we observed 92.44% of requests to originate from Canada and 99.83% of requests to originate from Europe from the corresponding bot services when inferring the geolocation from the IP address. Motivated by these results, we computed the geographical spread of requests based on both approaches and visualize them in Figure 10. The figure reveals a number of inconsistencies in geographical locations which also constitute spatial inconsistencies that can be leveraged to improve bot detection.
Takeaway 7: Bots deliver on promises of sending traffic from specific locations by altering their IP addresses or browser attributes or both. Any inconsistencies among the locations inferred using these avenues can help improve bot detection.
Figure 11: Temporal distribution of requests received on our honey site
6.3 Temporal inconsistencies across requests from the same device
@@
Figure 11 shows the temporal spread of requests received on our honey site over time. The plot shows the number of requests, the number of unique IP addresses, number of unique values for cookies set by our honey site and the number of unique FingerprintJS fingerprints seen per day.
@@
From the figure we see that even after 2 months, we receive requests with previously unseen fingerprints and IP addresses. More interestingly, the spikes in the plot correspond to days when we renewed our purchases. These spikes indicate that the bot services have access to a large number of devices with different device configurations that result in different browser attributes, and thus, different fingerprints. However, we suspect that bots have a fixed set of devices but manipulate browser attributes to create the illusion of sending requests from a large number of devices.
@@
Figure 12: Percentage of requests seen across different navigator platforms for the same Cookie. The diverse spread in this figure provides strong evidence of spoofing browser attributes such as the navigator platform. Unless the attribute was spoofed, all requests would have the same navigator platform.
@@
To provide evidence that bots manipulate their browser attributes, we inspect the value of the navigator object’s platform attribute on all requests bearing the most commonly seen Cookie. Whenever a device sends a request to our honey site, we store a large random number in a first-party Cookie if it has not been set. Thus, requests bearing the same value for this Cookie should originate from the same device. Since the platform property of the navigator object captures information about the type of processor on a given device, we expect to always extract the same value for that device unless the entity controlling the device has intentionally altered its value. In Figure 12, we see a wide distribution for the navigator’s platform property for the device identified as sending us the largest number of requests with the same Cookie. Differing values for browser attributes that cannot change for a given device constitutes a temporal inconsistency that can be used to improve bot detection.
@@
Takeaway 8: Bots change their browser attributes to create an illusion of sending requests from a large number of devices. However, recording differing values for browser attributes that cannot change for a given device constitutes temporal inconsistencies that provide an avenue to detect bots.
7 FP-Inconsistent
@@
Our measurements in Section 6 reveal that there exist inconsistencies in different browser attributes for a given request as well as multiple requests from the same device over a period of time. In this section we present our approach to use these inconsistencies to enhance bot detection. We define inconsistencies as feature values that are either incompatible (unable to occur simultaneously with other values) with other features within the same request or contradictory to the same feature across multiple requests. We categorize inconsistencies into two types: spatial and temporal.
@@
Spatial inconsistencies refer to feature values within a request that conflict or are incompatible with other feature values in that same request. Examples include differing locations inferred from an IP address and time zone, or implausible combinations, such as an iPhone without touch input support. Our takeaways in Section 6.1 and Section  6.2 show that evasive bots incur significant spatial inconsistencies across information provided by browser attributes and IP address.
@@
Temporal inconsistencies are feature values that are incompatible across different requests from the same user or users with the same IP address. Examples include significantly different time zones for requests from the same IP address and inconsistent device memory values for the same user. Our takeaway from Section 6.3 shows that evasive bots give rise to significant temporal inconsistencies in an attempt to evade detection by changing their attributes.
7.1 Identifying spatial inconsistencies
@@
Our methodology for detecting spatial inconsistencies relies on the understanding that real devices can only possess a limited number of hardware and software configurations. In contrast, bots, in their attempts to mimic real devices and evade detection, as described in Section 6.3, often modify these configurations. However, these alterations typically do not account for every possible source of device information (such as JavaScript APIs, User-Agent, etc.), leading to a proliferation of device configurations in our dataset. This is especially noticeable in devices such as iPhones or iPads that are commonly owned by real users and have the highest success rate in evading detection by bot services (as shown in Section 6.1). Consequently, the increased number of bots pretending to be popular devices results in a greater variety of configurations in our historical dataset.
@@
However, identifying such inconsistencies is challenging because analyzing all possible feature combinations is infeasible. To facilitate the analysis, we first categorize features into different groups based on the type of information each feature provides. For instance, features like Color Depth, Screen Resolution, and Touch Support are grouped because they all convey information about the device’s screen. Table VII in Appendix C shows the various groups used in our analysis, demonstrating how we categorize features to streamline the detection of inconsistencies.
@@
Figure 13: An example of excessive configurations of a device (iPhone) with the feature representing maximum touch points.
@@
Next, we analyze pairs of features within each category to identify spatial inconsistencies. For each pair, we rank the features based on the number of unique instances recorded in our dataset. For instance, in the pair UA Device and Maximum Touch Points, we sort UA Device in descending order by the number of unique Max Touch Points values associated with it. A genuine iPhone can only five simultaneous touch points. However, when bots imitate iPhones but report a different number of touch points, our dataset reveals an implausible number of unique combinations between UA Device and Maximum Touch Points. We start with the UA Device instance that has the highest number of unique combinations and identify cases where the combination of these two features is impossible. After identifying the inconsistent pair of feature values, we repeat the process with lower ranked unique combinations and other feature pairs. Algorithm 1 describes the process we use to identify spatial inconsistencies. This algorithm helps us identify the most frequently altered features and the spatial inconsistencies they produce. Table VI in Appendix B provides examples of such inconsistencies in our dataset.
Algorithm 1 Algorithm to Detect Spatial Inconsistencies
1:  Input: Feature categories $F$, Dataset containing requests $D$, Labels for requests $L$ (true if the request is from a bot, false for human)
2:  for all $f\in F$ do
3:     for all feature pairs $\{f_{a},f_{b}\}\subseteq f$ do
4:        Filter $D$ where $L$ is false, creating $D^{\prime}$
6:        Sort the tuples in increasing order of $nv_{f_{b}}$
7:        for all $(v_{f_{a}},nv_{f_{b}})$ in the sorted order do
8:           if the combination is inconsistent then
9:              Label all rows in $D$ containing $(v_{f_{a}},v_{f_{b}})$ as true
10:           end if
11:        end for
12:     end for
13:  end for
7.2 Identifying temporal inconsistencies
@@
Building upon our findings in Section 6.3, we utilize both the user identifier, set by our honey sites in each visiting user’s browser storage and IP address to identify temporal inconsistencies First, we use the user identifier to measure variance in immutable device features (e.g., number of CPU cores, device memory) across requests containing the same user identifier. If an incoming request increases the number of unique feature values associated with existing identifiers, we consider that request to be temporally inconsistent. For instance, if all previous requests from a user have a Hardware Concurrency value of 4 and a new request contains a value of 6, we label that request as temporally inconsistent.
@@
We also use a user’s IP address to identify temporal inconsistencies related to time zones and location. If an incoming request increases the number of unique time zones (measured as an offset from UTC) associated with that IP, we classify that request as temporally inconsistent.” Similarly, we also identify temporal inconsistencies in location information provided through the IP address and navigator.geolocation.
7.3 Accuracy improvement through inconsistency analysis
In this subsection, we describe our methodology to use temporal and spatial inconsistencies for enhancing the detection of bots that evade DataDome and BotD. To measure the improvement in accuracy from spatial inconsistencies, we translate the inconsistencies identified in Table VI into filter rules. These filter rules are then matched with each each request that evaded detection by DataDome or BotD. For temporal inconsistencies, we use the timestamp of each request to determine the order in which requests were made, applying filter rules to identify inconsistencies created by requests arriving later.
The results in Table III show that using rules generated through spatial and temporal inconsistency analysis can decrease the evasion of bots against BotD by 44.95% and against DataDome by 48.11%.
TABLE III: Comparison of improvement to BotD and DataDome accuracy with inconsistency analysis
BotD Datadome
None 47.07% 55.44%
Spatial 70.33% 76.04%
Temporal 48.09% 56.53%
Combined 70.86% 76.88%
@@ 7.3 Accuracy improvement through inconsistency analysis
TABLE IV: Improvement in DataDome and BotD’s detection rate on requests from each bot service when incorporating FP-Inconsistent

### v2

@@ 6. Inconsistency analysis
From our analysis in the previous section, we see ensuring certain values for certain fingerprint attributes helps bots evade detection. One way in which evasive bots could accomplish this would be to send requests from devices that would contain the desired values for attributes. For example, evasive bots could send requests from devices containing 4 CPU cores to ensure a value of 4 for hardwareConcurrency. Alternatively, evasive bots could alter browser APIs and device properties to present their desired values for fingerprint attributes (Liu et al., 2022). In this case, an evasive bot could alter the hardwareConcurrency attribute of the navigator object to return 4 on a device that may not have 4 CPU cores.
@@
In this section, we describe various inconsistencies in fingerprint attributes among the requests received on our honey site. These inconsistencies provide evidence of bots altering browser APIs since such inconsistencies are extremely unlikely to occur when using real devices. We use insights from these inconsistencies to develop FP-Inconsistent, our semi-automated technique to generate inconsistency rules to detect evasive bots (Section 7).
@@
Figure 6. Bar plot showing the top 4 device types (inferred from the User-Agent) that have the highest probability of evading DataDome.
6.1. Inconsistencies across fingerprint attributes
@@
Figure 6 shows the top 4 device types (inferred using the User-Agent property of the browser’s navigator object) that have the highest probability of evading DataDome among the requests recorded on our honey site. From the figure, we see that iPhones have the highest probability of evasion (around 50%). We now look at other fingerprint attributes to determine if evasive bots sent requests from real iPhones or if they altered the navigator object on their browser to have their devices appear as iPhones. Since iPhones have a fixed set of screen resolutions (12 resolutions (Eugene Belinski, [n. d.])), we inspect the spread of screen resolutions captured on requests from iPhones. Upon inspection, we found 83 unique screen resolutions from iPhones, out of which 42 were present among those requests from iPhones that evaded DataDome. We also find that 9 out of the top 10 screen resolutions that have the highest probability of evading DataDome among requests claiming to use iPhones do not exist in the real world. We visualize these probabilities in Figure 7. This provides strong evidence that bots alter browser APIs to show that they use iPhones rather than using actual iPhones.
From this evidence, we see that while bots alter browser APIs, it is difficult for them to ensure that all fingerprint attributes remain consistent with their alterations. Thus, inconsistencies across fingerprint attributes can be leveraged for bot detection since real users are unlikely to have such inconsistencies. In Section 7 we discuss our systematic, data-driven, semi-automatic approach to discover such inconsistencies to improve bot detection.
@@
Takeaway 4: While bots alter fingerprint attributes for evasion, they do not ensure that all attributes are consistent with their alteration. A particular value for a given attribute mapping to a large number of values for another attribute provides an avenue to discover inconsistencies.
Figure 7. Bar plot showing the top 10 screen resolutions among requests received from iPhones (inferred using the User-Agent) that have the highest probability of evasion against DataDome. 9 out of these 10 resolutions do not exist in the real world indicating an inconsistency that can be leveraged to detect bots.
@@
Figure 8. Plots showing a heatmap of the geographical location of requests inferred using the timezone attribute of the navigator object and the IP address. Different regions lighting up in the two heatmaps indicate that while bots alter the navigator object or IP address or both to change their geographical location, they do not ensure that the location inferred using both is consistent.
6.2. Inconsistencies across fingerprint attributes and IP addresses
@@
Some bot services advertised sending traffic from specific geographic regions (USA, Mexico, France, etc). Having this ability to send requests from specific regions suggests that the bot services are likely altering attributes that capture the geographical location of their devices. This alteration introduces potential inconsistencies if the bot services did not ensure that all attributes point to the same region.
@@
We analyzed requests from 4 different bot services who claimed to send requests from the United States, Canada, Europe, and France respectively. We first used MaxMind’s GeoLite2 database (MaxMind, 2024a) to extract the geolocation from the IP address of the requests from these services. We took a conservative approach when determining if the inferred geolocation matched the region advertised by the bot service. Concretely, we considered locations at the same UTC offset to be a match. For example, when analyzing requests from the bot service who advertised sending requests from France, we considered all requests whose geolocations mapped to any valid UTC offset that could overlap with France (such as Europe/Berlin) to also originate from France. With this approach, over 90% of requests from each of the 4 bot services matched the advertised geographical location.
@@
However, we observed significant differences when repeating the same analysis using the browser’s timezone API (Mozilla, [n. d.]) to infer location. We still used the same conservative approach and merely replaced the geolocation inferred from the IP address with the timezone. Only 76.52% of requests mapped to UTC offsets in Canada among the requests from the bot service that advertised traffic from Canada. More alarmingly, we observed that only 56% of requests mapped to UTC offsets in Europe among the requests from the bot service that advertised traffic from Europe. In contrast, we observed 92.44% of requests to originate from Canada and 99.83% of requests to originate from Europe from the corresponding bot services when inferring the geolocation from the IP address. Motivated by these results, we visualize the geographical spread of requests based on both approaches in Figure 8. The figure reveals a number of inconsistencies in geographical locations which also constitute inconsistencies for bot detection.
Takeaway 5: Bots alter their IP addresses, fingerprint attributes or both to fulfill promises of sending requests from specific locations.
Figure 9. Temporal distribution of traffic on our honey site.
6.3. Inconsistencies across time
@@
Figure 9 shows the temporal spread of requests received on our honey site over time. The plot shows the number of requests, the number of unique IP addresses, the number of unique values for Cookies set by our honey site, and the number of unique FingerprintJS fingerprints seen per day.
@@
From the figure, we see that even after 2 months, we receive requests with previously unseen fingerprints and IP addresses. More interestingly, the spikes in the plot correspond to the days when we renewed our purchases. These spikes indicate that the bot services could have access to a large number of devices with different device configurations that result in different browser attributes, and thus, different fingerprints. However, we suspect that they have a fixed set of devices but alter fingerprint attributes to create the illusion of sending requests from a large number of devices.
@@
Figure 10. Percentage of requests seen across different values of the platform attribute of the navigator object for the same Cookie (same device). The diverse spread of values provides strong evidence of bots altering the platform attribute since it cannot change otherwise for the same device.
@@
To provide evidence that bots alter their fingerprint attributes, we inspect the navigator object’s platform attribute on all requests that share the most commonly seen Cookie. Whenever a device sends a request to our honey site, we store a large random number in a first-party Cookie if it had not been set previously. Thus, requests bearing the same value for this Cookie should originate from the same device. Since the platform property of the navigator object captures information about the type of processor on a given device, it can never change for that device unless the entity controlling the device has intentionally altered the attribute. In Figure 10, we see a wide distribution for the navigator’s platform property for the device identified as sending us the largest number of requests with the same Cookie. Differing values for fingerprint attributes that cannot change for a given device constitutes a temporal inconsistency that can be used for bot detection.
@@
Takeaway 6: Bots alter fingerprint attributes to create an illusion of sending requests from a large number of devices. Recording differing values for fingerprint attributes that cannot change for a given device also constitute inconsistencies to detect bots.
7. FP-Inconsistent
@@
Our measurements in Section 6 reveal that there exist inconsistencies in different fingerprint attributes for a given request as well as multiple requests from the same device at different points in time. In this section, we present our approach to use these inconsistencies to enhance bot detection. We categorize inconsistencies into two types: spatial and temporal.
@@
Spatial inconsistencies refer to attribute values within a request that conflict or are incompatible with other attribute values in that same request. Examples include differing locations inferred from an IP address and time zone, or implausible combinations, such as an iPhone without touch input support. Our takeaways in Section 6.1 and Section  6.2 show that evasive bots incur significant spatial inconsistencies across information captured in their fingerprint attributes as well as IP addresses.
@@
Temporal inconsistencies are attribute values that are incompatible across different requests from the same user or users. Examples include significantly different time zones for requests from the same IP address and inconsistent device memory values for the same Cookie value. Our takeaway from Section 6.3 shows that evasive bots give rise to significant temporal inconsistencies by changing their attributes.
7.1. Identifying spatial inconsistencies
@@
Our methodology for detecting spatial inconsistencies relies on the understanding that real devices can only possess a limited number of hardware and software configurations. In contrast, bots, in their attempts to mimic real devices and evade detection, as described in Section 6.3, often modify these configurations. However, these alterations typically do not account for every possible source of device information (such as JavaScript APIs, User-Agent, etc.), leading to a proliferation of device configurations. This is especially noticeable in devices such as iPhones or iPads that are commonly owned by real users and have the highest success rate in evading detection (as shown in Section 6.1). Consequently, the increased number of bots pretending to be popular devices results in a greater variety of configurations in the dataset of requests obtained on our honey site.
@@
However, identifying such inconsistencies is challenging because analyzing all possible attribute combinations is infeasible. To facilitate the analysis, we first categorize attributes into different groups based on the type of information each attribute provides. For instance, attributes like Color Depth, Screen Resolution, and Touch Support are grouped because they all convey information about the device’s screen. Table 7 in Appendix F shows the various groups used in our analysis, demonstrating how we categorize attributes to streamline the detection of inconsistencies.
@@
Figure 11. An example of excessive configurations of a device (iPhone) with the fingerprint attribute representing maximum touch points.
@@
Next, we analyze pairs of attributes within each category to identify spatial inconsistencies. For each pair, we rank the attributes based on the number of unique instances recorded in our dataset. For example, in the pair UA Device and Maximum Touch Points, we sort UA Device in descending order by the number of unique Max Touch Points values associated with it. A genuine iPhone can only have five simultaneous touch points. However, when bots imitate iPhones but report a different number of touch points, our dataset reveals an implausible number of unique combinations between UA Device and Max Touch Points. We start with the UA Device instance that has the highest number of unique combinations and identify cases where the combination of these two attributes is impossible. After identifying the inconsistent pair of attribute values, we repeat the process with lower-ranked unique combinations and other attribute pairs. Appendix D defines our algorithm to identify spatial inconsistencies. This algorithm helps us identify the most frequently altered attributes and the spatial inconsistencies they produce. Table 6 in Appendix E provides examples of such inconsistencies in our dataset.
7.2. Identifying temporal inconsistencies
@@
Building upon our findings in Section 6.3, we utilize both the large random number identifier set by our honey sites in each visiting device’s browser storage (Cookie) and IP address to identify temporal inconsistencies. First, we use the Cookie identifier to measure variance in immutable device attributes (e.g., number of CPU cores, device memory) across requests containing the same identifier. If an incoming request increases the number of unique attribute values associated with previous identifiers, we consider that request to be temporally inconsistent. For instance, if all previous requests from a device have a Hardware Concurrency value of 4 and a new request contains a value of 6, we label that request as temporally inconsistent.
@@
We also use a user’s IP address to identify temporal inconsistencies related to time zones and location. If an incoming request increases the number of unique time zones (measured as an offset from UTC) associated with that IP, we classify that request as temporally inconsistent. Similarly, we also identify temporal inconsistencies in location information provided through the IP address and navigator.geolocation.
@@
Table 3. Improvement in DataDome and BotD’s detection rate on traffic from each bot service when incorporating FP-Inconsistent.
@@
7.3. Improved bot detection
In this section, we describe our methodology to use temporal and spatial inconsistencies to detect bots that evade DataDome and BotD. To measure the improvement in accuracy from spatial inconsistencies, we translate the inconsistencies identified in Table 6 into filter rules. These filter rules are then matched with each request that evaded detection by DataDome or BotD. For temporal inconsistencies, we use the timestamp of each request to determine the order in which requests were made, applying filter rules to identify inconsistencies created by requests arriving later.
The results in Table 4 show that using rules generated through spatial and temporal inconsistency analysis can decrease the evasion of bots against BotD by 44.95% and against DataDome by 48.11%. Table 3 shows the improvement in detection on requests obtained from each individual bot service. We evaluated the generalizability of our methodology by computing filter rules on 80% of the requests obtained on our honey site and evaluating them on the remaining 20%. This evaluation led to a meagre drop in detection accuracy of 0.42% for BotD and 0.23% for DataDome, thereby showing that FP-Inconsistent generalizes to unseen requests.
Table 4. Comparison of the improvement in DataDome and BotD’s detection accuracies resulting from different forms of inconsistency analysis.
DataDome BotD
None 55.44% 47.07%
Spatial 76.04% 70.33%
Temporal 56.53% 48.09%
Combined 76.88% 70.86%